OpenStudy (anonymous):

You are interested in buying a new home that costs $280,000. The mortgage company requires a down payment of 15% of the cost of the home. You will borrow the rest. The mortgage company offers you 2 options. Option 1 is a 15 year loan with an APR of 6.75% and option 2 is a 30 year loan with an APR of 5.55%. a. For each option, find the monthly mortgage payment, the total amount paid back and the total amount of interest paid. Write these in your Word document. b. Which option would you choose and why?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Step one is to find out what that down payment is.

OpenStudy (anonymous):

42,000

OpenStudy (anonymous):

residual amount is 238,000

OpenStudy (anonymous):

Alright, next: it's talking about monthly payments, so find out how many months 15 years is and how many months 30 years is.

OpenStudy (anonymous):

Also, you'll want to divide you APR (annual percentage rate) by 12 to get the monthly rate.

OpenStudy (anonymous):

180 months and 360 months

OpenStudy (anonymous):

what would be the next step im so lost

OpenStudy (anonymous):

0.005625 and 0.004625

OpenStudy (anonymous):

now what @cliffsedge

OpenStudy (anonymous):

Have you worked with exponential functions before? (Is this an algebra 2 class, or some other kind of course?)

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

college algebra

OpenStudy (anonymous):

i really needa get this done in the next 10 minutes =(

OpenStudy (anonymous):

Ok, so you have probably seen an exponential function before, something like \(\large y=a\cdot b^{kx}\)

OpenStudy (anonymous):

yes i have

OpenStudy (anonymous):

im staring so hard at my screen for an answer

OpenStudy (anonymous):

Great, so in this situation, the x in the exponent is the time (the number of months. The compound interest formula is \[A=P(1+R)^t\] A is the amount after t months, P is the 'principal' (the initial amount) R is the interest rate (in this case adjusted to be per month.)

OpenStudy (anonymous):

i did this but the number is soooo high i don't think its the answer

OpenStudy (anonymous):

If you're staring at the screen for an answer, you're looking in the wrong place. You need to be looking in your own brain for that. I can help you use what you already know to find it, but you have to do the work yourself.

OpenStudy (anonymous):

this is what i got

OpenStudy (anonymous):

42000*(1+.0675)^15 = 111883.887

OpenStudy (anonymous):

but isnt the formula P=(1+R/n)^n*T

OpenStudy (anonymous):

@cliffsedge

OpenStudy (anonymous):

It depends on what you are using for R. If you are using the APR as R, then yes, you need to divide by n (the number of payment periods) to adjust the annual rate to whatever other time period you are using (in this case, months, so n=12 periods per year).

OpenStudy (anonymous):

isnt it APR in this case

OpenStudy (anonymous):

I usually prefer to make the time period adjustments first and get a new, adjusted R, so the formula is a little easier to use.

OpenStudy (anonymous):

ok so whats the next step

OpenStudy (anonymous):

Yes, you may use R= APR, then you need to divide by n to adjust the rate, and multiply t by n to adjust the time. Like I said, I prefer to make the adjustments to the rate and the time first before putting things into the formula just to simplify the look of the formula. Either way will work out the same.

OpenStudy (anonymous):

ok sounds good

OpenStudy (anonymous):

I think (and I might be missing something because it's been a decade or so since I've worked in financial accounting, but I think all you need to do is use that formula to find out what the total amount will be after however many years, then divide by that many months to get the monthly payment.

OpenStudy (anonymous):

wait dont go anywhere let me do it real fast

OpenStudy (anonymous):

There are formulas coming to mind from my engineering economics class that make me think there is something else you need to do, but it's been nearly ten years since I've studied that in detail as well.

OpenStudy (anonymous):

i got 634008.6929

OpenStudy (anonymous):

I recommend you get a second opinion. I can check your algebra and make sure you're solving the equations right, but if there is a nuance to how the finances work that I'm missing, then I might lead you astray.

OpenStudy (anonymous):

That value seems a little high, but I'll check it; hold on....

OpenStudy (anonymous):

please helpppp

OpenStudy (anonymous):

For the 15 year, 6.75% loan, I got a total amount of $653,232.85 So, you answer is in the right area . . . how did you get it?

OpenStudy (anonymous):

i used the formula u gave me

OpenStudy (anonymous):

wait but ok u have to pay that much off to fullfill repayment? from a 238,000 loan?

OpenStudy (anonymous):

At nearly 7% per year for 15 years, ... yeah, that interest adds up. However, from my experience with this sort of thing, I know that that only gives an estimate because if you make a payment each month, then the interest rate is applied to the balance that you owe after making the payment.

OpenStudy (anonymous):

ok so u got that answer for the 15 years? can u tell me how u got that whic formula u used cuz i cant seem to come up with that answer

OpenStudy (anonymous):

Sure. The 6.75% APR is 0.005625 on a monthly basis. 15 years = 180 months. \(\large A=238,000(1.005625)^{180}\)

OpenStudy (anonymous):

waitttt where did u get that .005625 from

OpenStudy (anonymous):

it takes forever for a reply

OpenStudy (anonymous):

omfg i coulda been done by now

OpenStudy (anonymous):

this is one friggin problem omfg..........

OpenStudy (anonymous):

hero do u kno how to do this

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@agent_sniffles

hero (hero):

Okay, I'll post the steps

hero (hero):

Down Payment = .15(280,000) = 42,000 Borrow Amount: = 280,000 - 42000 = 238,000 Option 1: A = 238000(1.0675)^15 A = 634,009 Option 2: A = 238000(1.0555)^30 A = 1, 203,160 I would choose option 1 because it costs less. Better to get it over with and pay less than to pay more over a longer period.

OpenStudy (anonymous):

hmmmmmmmmmmm i got exactly that

OpenStudy (anonymous):

but what about everything else? how do i get the mortgage payment?

OpenStudy (anonymous):

For each option, find the monthly mortgage payment, the total amount paid back and the total amount of interest paid

OpenStudy (anonymous):

monthly mortgage payment do i just divide by how many months 15 years equals to?

hero (hero):

I didn't compute monthly mortgage amount. Looks like cliffsedge got that covered.

OpenStudy (anonymous):

i dont understand how he does it please explain to me hero my hero

OpenStudy (anonymous):

Does there appear to be anything I missed for getting the monthly amount, @Hero ? The last course I took which covered this was Engineering Economics, and it was quite complicated, so I might be over-thinking things in that mode. Since then, I've only ever used exponentials to pay off my credit cards, so maybe mortgages are different?

hero (hero):

He converted APR and number of years to monthly payments. 15 years = 180 months 6.75/12 = .5675

hero (hero):

Which is the correct approach. If you use that you get: A = 238000(1.005675)^180 A = 653,233

OpenStudy (anonymous):

ohhhhhhh ok i see now

OpenStudy (anonymous):

but what is that number is that the total amount paid back that you just found?

hero (hero):

I'm not really sure how to find the monthly payment. But there are calculators that you can use.

OpenStudy (anonymous):

wait but what did you just figure out though how does that differ from the first one of 634,009

hero (hero):

2106.08 I get that for monthly payments using the calculator.

OpenStudy (anonymous):

can you tell me how you got that?

OpenStudy (anonymous):

^ Heh, I used to work for bankrate.com for like three years. I was an accountant there. I've forgotten how to do all this stuff since then. :">

OpenStudy (anonymous):

Sorry I couldn't be more helpful.

OpenStudy (anonymous):

how do u forget if u worked doing this kind of stuff

OpenStudy (anonymous):

i got 2477.75

OpenStudy (anonymous):

wait but what is this though 653,233 how does that differ from the first one of 634,009

hero (hero):

Don't pay attention to the 634,009. It was computed using the annual percentage rate rather than monthly percentage rate.

OpenStudy (anonymous):

@yomamabf "how do u forget if u worked doing this kind of stuff " It was over 15 years ago, and I've been doing physics and engineering since then, so my mind has been immersed in other equations since. Sorry, but the human brain can only remember so many things over time. I guess I should have studied more. :-p

hero (hero):

Yeah, it is easier to document course work you have done over a period of time rather than try to remember it all. That way, you can refer to it later as reference material.

OpenStudy (anonymous):

@hero um hello?wait but what is this though 653,233 how does that differ from the first one of 634,009

OpenStudy (anonymous):

help me satelliteeeeeeeeeeee

OpenStudy (anonymous):

i hate to throw cold water on this, but i don't think those are the right payment formulas

OpenStudy (anonymous):

oh sorry, i missed @hero link let me shut up

OpenStudy (anonymous):

you have to borrow 85% of 280,000 which is 238,000 use the mortgage calculator link hero sent

OpenStudy (anonymous):

15 years at 6.75 gives a monthly payment of $2106.08

OpenStudy (anonymous):

30 years at 5.55 gives a monthly payment of $1358.81

OpenStudy (anonymous):

total payment over 15 years is \(180\times 2106.08= 379,094.40\) and total payment for 30 years is \(360\times 1358.81=489,171.60\)

OpenStudy (anonymous):

you pay a lot more over the course of the loan with the 30 year mortgage, on the other hand maybe you cannot afford such high monthly payments that the 15 year mortgage has besides, maybe you will move in 5 years

OpenStudy (anonymous):

thank you satellite73 god of math! *bows*

hero (hero):

Yeah, but @satellite didn't explain how to find the monthly payment without using the link.

OpenStudy (anonymous):

way to steal his shine kanye west

OpenStudy (anonymous):

i can do that if you like, but it is boring

OpenStudy (anonymous):

\[pmt=\frac{P\times \frac{apr}{12}}{1-(1+\frac{apr}{12})^{-12Y}}\]

hero (hero):

I wasn't trying to steal anyone's shine

hero (hero):

I was just stating the obvious.

OpenStudy (anonymous):

assuming monthly payments borrowing P dollars at apr written as a decimal for Y years

hero (hero):

@satellite, Did you derive that formula yourself or did you find it in a book?